Peru Industries began operations on January 1, 2020.

Problem 8-6A Recording accounts receivable transactions and bad debt adjustments LO1, 2, 3

Peru Industries began operations on January 1, 2020. During the next two years, the company completed a number of transactions involving credit sales, accounts receivable collections, and bad debts (assume a perpetual inventory system). These transactions are summarized as follows:

2020

  1. Sold merchandise on credit for $2,310,000, terms n/30 (COGS = $1,276,000).

  2. Wrote off uncollectible accounts receivable in the amount of $35,200.

  3. Received cash of $1,378,000 in payment of outstanding accounts receivable.

  4. In adjusting the accounts on December 31, concluded that 1.5% of the outstanding accounts receivable would become uncollectible.


2021

  1. Sold merchandise on credit for $3,024,000, terms n/30 (COGS = $1,646,000).

  2. Wrote off uncollectible accounts receivable in the amount of $54,800.

  3. Received cash of $2,282,000 in payment of outstanding accounts receivable.

  4. In adjusting the accounts on December 31, concluded that 1.5% of the outstanding accounts receivable would become uncollectible.

Company uses the allowance method to account for uncollectible.


Required:
Prepare journal entries to record Peru’s 2020 and 2021 summarized transactions and the adjusting entries to record bad debt expense at the end of each year. (Round your intermediate calculations and final answers to nearest whole dollar.)
Answer #1

2020

2020
Account Titles Debit Credit
Accounts Receivable $   23,10,000
       Sales Revenue     $   23,10,000
(Sale of goods on account)
Cost of Goods Sold $   12,76,000
       Inventory $   12,76,000
(Cost of goods sold)
Allowance for Doubtful debts $        35,200
       Accounts Receivable $        35,200
(Accounts receivable written off)
Cash $   13,78,000
      Accounts Receivable $   13,78,000
(collected cash on accounts)
Bad Debt Expense $        48,652 =(2310000-35200-1378000)*1.5%+35200
      Allowance for Doubtful debts $        48,652
(Bad Debt expense for year)


Accounts Receivable Balance = 2310000-35200-1378000 = $896800

2021

Account Titles Debit Credit
Accounts Receivable $   30,24,000
       Sales Revenue     $   30,24,000
(Sale of goods on account)
Cost of Goods Sold $   16,46,000
       Inventory $   16,46,000
(Cost of goods sold)
Allowance for Doubtful debts $        54,800
       Accounts Receivable $        54,800
(Accounts receivable written off)
Cash $   22,82,000
      Accounts Receivable $   22,82,000
(collected cash on accounts)
Bad Debt Expense $        29,908 =(896800+3024000-54800-2282000)*1.5%-48652+54800
      Allowance for Doubtful debts $        29,908
(Bad Debt expense for year)